OpenStudy (mrhirohito):

Simplify the expression. A. 52 B. 48 C. 16 D. 22

OpenStudy (shelbywyatt):

what expression

OpenStudy (shelbywyatt):

without blue please

OpenStudy (juliannegirltomboy):

first 15x5

OpenStudy (juliannegirltomboy):

=75

OpenStudy (juliannegirltomboy):

then 75+-24=51

OpenStudy (juliannegirltomboy):

51+18=69

OpenStudy (g3n3sis08):

its 16.

OpenStudy (shelbywyatt):

15*5=75 75-24=51 51-18=33 33*2/3=x

OpenStudy (shelbywyatt):

33/1 * 2/3 simplied to 11/1* 2/1

OpenStudy (shelbywyatt):

which 11*2= 22 1*1=1 22/1

OpenStudy (shelbywyatt):

22/1=22

OpenStudy (g3n3sis08):

15x5=75 75-24=51 51x2/3=34 34-18=16 follow the brackets

OpenStudy (g3n3sis08):

2/3 is multiplied by 51 not 18

OpenStudy (shelbywyatt):

so?

OpenStudy (shelbywyatt):

the answer is ?

OpenStudy (g3n3sis08):

2/3x((75)-24) + ( -18) 2/3x(51) + (-18) 34+(-18) 34-18=16 answer is 16 or C

OpenStudy (g3n3sis08):

2/3x[(75)-24] + ( -18) *